OpenStudy (anonymous):

Four milliliters of ink A is added to 10 milliliters of ink B, forming a mixture of inks. Ink A contains 10% blue pigment. Ink B contains an unknown percentage of blue pigment. The mixture of the two inks contains 35% blue pigment. What percentage of blue pigment is in the mixture of the two inks? 45% 49% 60% 75%

OpenStudy (anonymous):

Ans is 45% Blue pigment in ink A = 4ml*(10/100) Blue pigment in ink B = 10ml*(x/100) Blue pigment in mixture = 14ml*(35/100)

OpenStudy (anonymous):

The amount of Blue pigment is fixed so 4*(10/100)+10*(x/100) = 14*(35/100) x = 45